# Quantifying the noise sensitivity of the Wasserstein metric for images

### v1

@@ Abstract
Wasserstein metrics are increasingly being used as similarity scores for images treated as discrete measures on a grid, yet their behavior under noise remains poorly understood. In this work, we consider the sensitivity of the signed Wasserstein distance with respect to pixel-wise additive noise and derive non-asymptotic upper bounds. Among other results, we prove that the error in the signed 2-Wasserstein distance scales with the square root of the noise standard deviation, whereas the Euclidean norm scales linearly. We present experiments that support our theoretical findings and point to a peculiar phenomenon where increasing the level of noise can decrease the Wasserstein distance. A case study on cryo-electron microscopy images demonstrates that the Wasserstein metric can preserve the geometric structure even when the Euclidean metric fails to do so.
@@ 1 Introduction
Optimal Transport (OT) provides a principled way to measure the distance between probability measures, capturing not only pointwise differences but also the underlying geometry of the data. Recent advances in computational approximation methods (Cuturi, 2013; Schmitzer, 2019) contributed greatly to the rising popularity of optimal transport across many domains, such as computer vision (Feydy et al., 2021), domain adaptation (Courty et al., 2017), and others. In imaging applications, the Wasserstein metric can be used to measure similarity by treating images as discrete measures on a grid, and assigning a point mass to every pixel, proportional to its value. One field where this approach is gaining popularity is in single-particle cryo-electron microscopy (cryo-EM), a domain characterized by extremely high noise levels, where OT-based methods have been successfully applied to fundamental tasks, including the alignment of 3D density maps (Riahi et al., 2022; Singer & Yang, 2024), the clustering 2D tomographic projections (Rao et al., 2020), and the rotational alignment of tomographic projections with heterogeneity (Shi et al., 2025). We believe that a major driver for this adoption is that, empirically, the Wasserstein metric appears more robust to noise than the standard Euclidean norm.
In machine learning, OT-based metrics have inspired approaches that exhibit improved stability in training deep neural networks, most notably Wasserstein generative adversarial networks (WGAN) (Arjovsky et al., 2017) and Wasserstein autoencoders (WAE) (Tolstikhin et al., 2017). In these methods, the exact Wasserstein distance is not calculated, and is replaced by approximations, such as adversarially learned critics in WGANs or penalized divergences in WAEs, that serve as proxies for the true metric. The suggested explanation for their success is that a noise-robust metric creates a smoother loss landscape, which prevents a model from overfitting to irrelevant perturbations in the training data (Li et al., 2018). Despite these empirical successes, the theoretical understanding of the noise sensitivity of OT-based metrics remains limited. This paper aims to be a first step to bridge that gap.
@@
On the theoretical side, we provide quantitative bounds relating the signed Wasserstein cost (see equation 3) between noise-corrupted images and the signed Wasserstein cost between the clean images. Focusing on a Gaussian noise model with fixed mass and pixel-wise standard deviation proportional to $\sigma$, we show that the signed $p$-Wasserstein metric between a noise-corrupted $n\times n$ picture and its clean counterpart gives rise to an error term that scales like $(n\sigma)^{1/p}$ up to log terms, see Theorem 3. For the 1-Wasserstein distance, considering a similar noise model, Theorem 4 establishes that the distance between two noisy pictures deviate at most by an order $\sigma n\log_{2}n$ from the distance between the clean ones. Theorem 5 gives a bound for the case of two different measures and $p\geq 1$. We complement our theoretical results with simulations in Section 4, showcasing the properties of the signed Wasserstein distance in a variety of cases.
Figure 1: Distance ratios of $L^{2},W_{1}$ and $W_{2}$ on a pair of noisy images as a function of the noise level. $L^{2}$ diverges first, followed by $W_{1}$ and lastly, $W_{2}$ departs from the original distance between the images, exhibiting more noise robustness. Above each marker we show the pair of images that were compared using all 3 metrics. See Section 4.2 for more details.
2 Wasserstein over noisy and signed measures
@@
Under mild conditions—see Santambrogio (2015, Theorem 1.39), the Wasserstein distance admits a dual formulation, i.e.,
@@
Some image modalities (such as cryo-EM) naturally involve negative pixels, but even in modalities where all pixels are positive, once pixel-wise noise is introduced, negative pixels may appear. Since we identify pixel values with point masses, to study the effect of pixel-wise noise we first need to explain how the Wasserstein metrics can be extended to support negative masses.
The problem of generalizing OT to signed measures is not new; Mainini (2012) proposed to compute a Wasserstein-like distance between $\mu$ and $\nu$ by first constructing
@@
where we assume that the total mass of $\mu$ and $\nu$ are equal and $\mu_{+}$ ($\mu_{-}$, resp.) denotes the positive (resp. negative) part of $\mu$. Mainini (2012) then introduced the signed Wasserstein cost,
@@
which the author denoted by $\mathbb{W}_{p}(\mu,\nu)$. We note that $W^{\pm}_{p}$ is a metric for $p=1$ but not for $p>1$, since it does not satisfy the triangle inequality (Mainini, 2012, Proposition 3.4). The fact that the 1-Wasserstein metric combines nicely with the positive and negative parts can be deduced from equation equation 2. The absence of triangle inequality might further lead to surprising behaviors as seen in Figure 4.
Other approaches for generalizing the Wasserstein distance to signed measures were explored by Engquist et al. (2016). Further, the usefulness of signed Wasserstein costs as defined above starts to be acknowledged in the statistics literature as the recent preprint by Groppe et al. (2025) suggests.
@@
$\mu_{\epsilon}:=\mu+{\epsilon}_{\mu}\quad\text{ and }\quad\nu_{\epsilon}:=\nu+{\epsilon}_{\nu}$ (5)
@@
$S_{\mu_{\epsilon},\nu_{\epsilon}}:=(\mu_{\epsilon})_{+}+(\nu_{\epsilon})_{-}\quad\text{ as well as }\quad T_{\mu_{\epsilon},\nu_{\epsilon}}:=(\nu_{\epsilon})_{+}+(\mu_{\epsilon})_{-}.$ (6)
@@
$C_{S}:=\sum_{x\in G_{n}}(\mu_{\epsilon})_{+}(x)+(\nu_{\epsilon})_{-}(x)\quad\text{ and }\quad C_{T}:=\sum_{x\in G_{n}}(\nu_{\epsilon})_{+}(x)+(\mu_{\epsilon})_{-}(x).$ (7)
Standardizing $S_{\mu_{\epsilon},\nu_{\epsilon}}$ by $C_{S}$ and $T_{\mu_{\epsilon},\nu_{\epsilon}}$ by $C_{T}$ is necessary to ensure that both measures have the same (unit) mass in the case where $\sum_{x\in G_{n}}\mu_{\epsilon}(x)\neq\sum_{x\in G_{n}}\nu_{\epsilon}(x)$. In the sequel, we will use the notation
$\bar{S}_{\mu_{\epsilon},\nu_{\epsilon}}:=\frac{S_{\mu_{\epsilon},\nu_{\epsilon}}}{C_{S}}\quad\text{ and }\quad\bar{T}_{\mu_{\epsilon},\nu_{\epsilon}}:=\frac{T_{\mu_{\epsilon},\nu_{\epsilon}}}{C_{T}}.$ (8)
We aim at understanding the relationship between $W^{\pm}_{p}(\mu,\nu)$ and $W_{p}(\bar{S}_{\mu_{\epsilon},\nu_{\epsilon}},\bar{T}_{\mu_{\epsilon},\nu_{\epsilon}})$. To put our analysis into context, consider the standard squared $L^{2}$ distance, a common metric for image comparison. In the presence of additive Gaussian noise with variance $\sigma^{2}$, the expected squared $L^{2}$ distance between a signal and its noisy version has a simple, direct relationship: it is exactly $n\sigma$. This metric, however, is local and insensitive to the underlying geometric structure of the signal. In contrast, the (signed) Wasserstein cost is claimed to capture this geometry, but its behavior under noise is far more complex to characterize. This paper aims to bridge that gap by providing a theoretical and empirical analysis of its robustness.
@@
To get sharp estimates on the Wasserstein distance, the following proposition is particularly useful. This is Proposition 1 of Weed & Bach (2019), but on a domain with an arbitrary diameter (their formulation assumed $\operatorname{diam}(S)=1$). The bound is based on the construction of a coupling at various scales, managing the mass imbalance in subdomains. This construction yields sharp rates in a variety of cases.
@@ 3 Theoretical Contributions
$W_{p}^{p}(\bar{S}_{\mu_{\epsilon},\nu_{\epsilon}},\bar{T}_{\mu_{\epsilon},\nu_{\epsilon}})=\sup_{f}\langle f,\bar{S}_{\mu_{\epsilon},\nu_{\epsilon}}\rangle+\langle f^{\mathsf{d}},\bar{T}_{\mu_{\epsilon},\nu_{\epsilon}}\rangle$ (10)
$=\sup_{f}\frac{\left\langle f,S_{\mu_{\epsilon},\nu_{\epsilon}}\right\rangle+\left\langle f^{\mathsf{d}},T_{\mu_{\epsilon},\nu_{\epsilon}}\right\rangle}{\sum_{x\in G_{n}}S_{\mu_{\epsilon},\nu_{\epsilon}}(x)}+\left(\tfrac{1}{\sum_{x\in G_{n}}T_{\mu_{\epsilon},\nu_{\epsilon}}(x)}-\tfrac{1}{\sum_{x\in G_{n}}S_{\mu_{\epsilon},\nu_{\epsilon}}(x)}\right)\left\langle f^{\mathsf{d}},T_{\mu_{\epsilon},\nu_{\epsilon}}\right\rangle.$ (11)
This decomposition shows that the optimal dual function must balance two objectives at the same time: the first one is the transport problem, and the second can be interpreted as a mass imbalance penalization.
In the case of i.i.d. Gaussian noise, the result above can be refined and yields the following theorem.
@@
Consider two $n\times n$ images $\mu$ and $\nu$ having at least $\lambda n^{2}$, $\lambda\in(0,1]$ nonzero pixels. Assume that ${\epsilon}_{\mu},{\epsilon}_{\nu}$ are $\mathcal{N}(0_{n^{2}},\sigma^{2}I_{n^{2}})$. Recall the definition of $\bar{S}_{\mu_{\epsilon},\nu_{\epsilon}},\bar{T}_{\mu_{\epsilon},\nu_{\epsilon}}$ in equation 8. Then,
$W_{1}^{\pm}(\bar{S}_{\mu_{\epsilon},\nu_{\epsilon}},\bar{T}_{\mu_{\epsilon},\nu_{\epsilon}})=\frac{1}{\sum_{x\in G_{n}}S_{\mu_{\epsilon},\nu_{\epsilon}}(x)}\sup_{f\in\operatorname{Lip}_{1}}\left\langle f,S_{\mu_{\epsilon},\nu_{\epsilon}}-T_{\mu_{\epsilon},\nu_{\epsilon}}\left(1+\mathrm{O}_{p}\left(\frac{\sigma}{n}\right)\right)\right\rangle.$ (12)
@@
$\sum_{x\in G_{n}}S_{\mu_{\epsilon},\nu_{\epsilon}}(x)-T_{\mu_{\epsilon},\nu_{\epsilon}}(x)=\mathrm{O}_{p}(\sigma n),$ (13)
@@
In general, one can hope that the ratio $\sigma/n$ is small, so that the result suggests that understanding the quantity $\sup_{f}\left\langle f,S_{\mu_{\epsilon},\nu_{\epsilon}}\right\rangle+\left\langle f^{\mathsf{d}},T_{\mu_{\epsilon},\nu_{\epsilon}}\right\rangle$ under a suitable choice of noise is a first step to take towards completely characterizing the impact of the noise.
@@ 3.2 Noise model
The previous section invites us to consider a noise model for which it is not necessary to rescale the measures. To this end, we will consider slightly correlated Gaussian noise.
@@
This last property is the crucial point enabling to focus on the impact of the noise, while setting aside the potential questions pertaining to rescaling the measures whose behavior was captured in Theorem 1.

### v2

@@ Abstract
Wasserstein metrics are increasingly being used as similarity scores for images treated as discrete measures on a grid, yet their behavior under noise remains poorly understood. In this work, we consider the sensitivity of the signed Wasserstein distance with respect to pixel-wise additive noise and derive non-asymptotic upper bounds. Among other results, we prove that the error in the signed 2-Wasserstein distance scales with the square root of the noise standard deviation, whereas the Euclidean norm scales linearly. We present experiments that support our theoretical findings [^1] Code available at Github Repo and point to a peculiar phenomenon where increasing the level of noise can decrease the Wasserstein distance. A case study on cryo-electron microscopy images demonstrates that the Wasserstein metric can preserve the geometric structure even when the Euclidean metric fails to do so.
@@ 1 Introduction
Optimal Transport (OT) provides a principled way to measure the distance between probability measures, capturing not only pointwise differences but also the underlying geometry of the data. Recent advances in computational approximation methods (Cuturi2013; Schmitzer2019) contributed greatly to the rising popularity of optimal transport across many domains, such as computer vision (FeydyRoussillonTrouveGori2021), domain adaptation (CourtyFlamaryTuiaRakotomamonjy2017), and others. In imaging applications, the Wasserstein metric can be used to measure similarity by treating images as discrete measures on a grid, and assigning a point mass to every pixel, proportional to its value. One field where this approach is gaining popularity is in single-particle cryo-electron microscopy (cryo-EM), a domain characterized by extremely high noise levels, where OT-based methods have been successfully applied to fundamental tasks, including the alignment of 3D density maps (RiahiWoollardPoitevinCondonDuc2022; SingerYang2024), the clustering 2D tomographic projections (RaoMoscovichSinger2020), and the rotational alignment of tomographic projections with heterogeneity (ShiSingerVerbeke2025). We believe that a major driver for this adoption is that, empirically, the Wasserstein metric appears more robust to noise than the standard Euclidean norm.
Related work.
In generative modeling, OT-based metrics have inspired methods such as Wasserstein GAN (ArjovskyChintalaBottou2017) and Wasserstein autoencoders (TolstikhinBousquetGellySchoelkopf2017) and flow matching (LipmanChenBen-HamuNickelLe2022; AlbergoVanden-Eijnden2022; LiuGongLiu2022). The latter in particular has strong connections to OT in its dynamic formulation. Building upon this, recent variants of flow matching incorporate OT solvers into the training process (TongFatrasMalkinHuguetZhangEtAl2023; ChemseddineHagemannSteidlWald2025; ZhangMousavi-HosseiniKleinCuturi2025; Mousavi-HosseiniZhangKleinCuturi2025). While our work does not target these models specifically, we believe that a better understanding of the noise robustness of optimal transport procedures is needed to understand why modern generative models work so well.
Many variants of OT such as partial optimal transport (ChapelAlayaGasso2020; RaghvendraShirzadianZhang2024a) and unbalanced optimal transport (BenamouCarlierCuturiNennaPeyre2015; ChizatPeyreSchmitzerVialard2018) have been proposed to address mass imbalance. While not strictly comparable to the our work, these works cover a different notion of robustness.
@@
On the theoretical side, we provide quantitative bounds relating the signed Wasserstein cost (see equation 3) between noise-corrupted images and the signed Wasserstein cost between the clean images. Focusing on a Gaussian noise model with fixed mass and pixel-wise standard deviation proportional to $\sigma$, we show that the signed $p$-Wasserstein metric between a noise-corrupted $n\times n$ picture and its clean counterpart gives rise to an error term that scales like $(n\sigma)^{1/p}$, see Theorem 3. For the 1-Wasserstein distance, considering a similar noise model, Theorem 4 establishes that the distance between two noisy pictures deviate at most by an order $\sigma n\log_{2}n$ from the distance between the clean ones. Theorem 5 gives a bound for the case of two different measures and $p\geq 1$. We complement our theoretical results with simulations in Section 4, showcasing the properties of the signed Wasserstein distance in a variety of cases.
Figure 1: Distance ratios of $L^{2},W_{1}$ and $W_{2}$ on a pair of noised images as a function of the noise level. $L^{2}$ diverges first, followed by $W_{1}$ and lastly, $W_{2}$ departs from the original distance between the images, exhibiting more noise robustness. Above each marker we show the pair of images that were compared using all 3 metrics. See Section 4.2 for more details.
2 Wasserstein over noised and signed measures
@@
Under mild conditions, see Santambrogio2015, the Wasserstein distance admits a dual formulation, i.e.,
@@
Some image modalities (such as cryo-EM) naturally involve negative pixels, but even in modalities where all pixels are positive, once pixel-wise noise is introduced, negative pixels may appear. Since we identify pixel values with point masses, to study the effect of pixel-wise noise we first need to explain how the Wasserstein metrics can be extended to support negative masses. The problem of generalizing OT to signed measures is not new; Mainini2012 proposed to compute a Wasserstein-like distance between $\mu$ and $\nu$ by first constructing
@@
where we assume that the total mass of $\mu$ and $\nu$ are equal and $\mu_{+}$ ($\mu_{-}$, resp.) denotes the positive (resp. negative) part of $\mu$. Mainini2012 then introduced the signed Wasserstein cost,
@@
which the author denoted by $\mathbb{W}_{p}(\mu,\nu)$. We note that $W^{\pm}_{p}$ is a metric for $p=1$ but not for $p>1$, since it does not satisfy the triangle inequality (Mainini2012, Proposition 3.4). The fact that the 1-Wasserstein metric combines nicely with the positive and negative parts can be deduced from equation equation 2. The absence of triangle inequality might further lead to surprising behaviors as seen in Figure 4.
Other approaches for generalizing the Wasserstein distance to signed measures were explored by EngquistFroeseYang2016. Further, the usefulness of signed Wasserstein costs as defined above starts to be acknowledged in the statistics literature as the recent preprint by GroppeNiemollerHundrieserVentzkeBlobEtAl2025 suggests.
@@
$\mu_{\varepsilon}:=\mu+\varepsilon_{\mu}\quad\text{ and }\quad\nu_{\varepsilon}:=\nu+\varepsilon_{\nu}$ (5)
@@
$S_{\mu_{\varepsilon},\nu_{\varepsilon}}:=(\mu_{\varepsilon})_{+}+(\nu_{\varepsilon})_{-}\quad\text{ as well as }\quad T_{\mu_{\varepsilon},\nu_{\varepsilon}}:=(\nu_{\varepsilon})_{+}+(\mu_{\varepsilon})_{-}.$ (6)
@@
$C_{S}:=\sum_{x\in G_{n}}(\mu_{\varepsilon})_{+}(x)+(\nu_{\varepsilon})_{-}(x)\quad\text{ and }\quad C_{T}:=\sum_{x\in G_{n}}(\nu_{\varepsilon})_{+}(x)+(\mu_{\varepsilon})_{-}(x).$ (7)
Standardizing $S_{\mu_{\varepsilon},\nu_{\varepsilon}}$ by $C_{S}$ and $T_{\mu_{\varepsilon},\nu_{\varepsilon}}$ by $C_{T}$ is necessary to ensure that both measures have the same (unit) mass in the case where $\sum_{x\in G_{n}}\mu_{\varepsilon}(x)\neq\sum_{x\in G_{n}}\nu_{\varepsilon}(x)$. In the sequel, we will use the notation
$\bar{S}_{\mu_{\varepsilon},\nu_{\varepsilon}}:=\frac{S_{\mu_{\varepsilon},\nu_{\varepsilon}}}{C_{S}}\quad\text{ and }\quad\bar{T}_{\mu_{\varepsilon},\nu_{\varepsilon}}:=\frac{T_{\mu_{\varepsilon},\nu_{\varepsilon}}}{C_{T}}.$ (8)
We aim at understanding the relationship between $W^{\pm}_{p}(\mu,\nu)$ and $W_{p}(\bar{S}_{\mu_{\varepsilon},\nu_{\varepsilon}},\bar{T}_{\mu_{\varepsilon},\nu_{\varepsilon}})$. To put our analysis into context, consider the standard squared $L^{2}$ distance, a common metric for image comparison. In the presence of additive Gaussian noise with variance $\sigma^{2}$, the expected squared $L^{2}$ distance between a signal and its noisy version has a simple, direct relationship: it is exactly $n\sigma$. This metric, however, is local and insensitive to the underlying geometric structure of the signal. In contrast, the (signed) Wasserstein cost is claimed to capture this geometry, but its behavior under noise is far more complex to characterize. This paper aims to bridge that gap by providing a theoretical and empirical analysis of its robustness.
@@
To get sharp estimates on the Wasserstein distance, the following proposition is particularly useful. This is Proposition 1 of WeedBach2019, but on a domain with an arbitrary diameter (their formulation assumed $\operatorname{diam}(S)=1$). The bound is based on the construction of a coupling at various scales, managing the mass imbalance in subdomains. This construction yields sharp rates in a variety of cases.
@@ 3 Theoretical Contributions
$W_{p}^{p}(\bar{S}_{\mu_{\varepsilon},\nu_{\varepsilon}},\bar{T}_{\mu_{\varepsilon},\nu_{\varepsilon}})=\sup_{f}\langle f,\bar{S}_{\mu_{\varepsilon},\nu_{\varepsilon}}\rangle+\langle f^{\mathsf{d}},\bar{T}_{\mu_{\varepsilon},\nu_{\varepsilon}}\rangle$ (10)
$=\sup_{f}\frac{\left\langle f,S_{\mu_{\varepsilon},\nu_{\varepsilon}}\right\rangle+\left\langle f^{\mathsf{d}},T_{\mu_{\varepsilon},\nu_{\varepsilon}}\right\rangle}{\sum_{x\in G_{n}}S_{\mu_{\varepsilon},\nu_{\varepsilon}}(x)}+\left(\tfrac{1}{\sum_{x\in G_{n}}T_{\mu_{\varepsilon},\nu_{\varepsilon}}(x)}-\tfrac{1}{\sum_{x\in G_{n}}S_{\mu_{\varepsilon},\nu_{\varepsilon}}(x)}\right)\left\langle f^{\mathsf{d}},T_{\mu_{\varepsilon},\nu_{\varepsilon}}\right\rangle.$ (11)
This decomposition shows that the optimal dual function must balance two objectives at the same time: the first one is the transport problem, and the second can be interpreted as a mass imbalance penalization. In the case of i.i.d. Gaussian noise, the result above can be refined to yield,
@@
Consider two $n\times n$ images $\mu$ and $\nu$ having at least $\lambda n^{2}$, $\lambda\in(0,1]$ nonzero pixels. Assume that $\varepsilon_{\mu},\varepsilon_{\nu}$ are $\mathcal{N}(0_{n^{2}},\sigma^{2}I_{n^{2}})$. Recall the definition of $\bar{S}_{\mu_{\varepsilon},\nu_{\varepsilon}},\bar{T}_{\mu_{\varepsilon},\nu_{\varepsilon}}$ in equation 8. Then,
$W_{1}^{\pm}(\bar{S}_{\mu_{\varepsilon},\nu_{\varepsilon}},\bar{T}_{\mu_{\varepsilon},\nu_{\varepsilon}})=\frac{1}{\sum_{x\in G_{n}}S_{\mu_{\varepsilon},\nu_{\varepsilon}}(x)}\sup_{f\in\operatorname{Lip}_{1}}\left\langle f,S_{\mu_{\varepsilon},\nu_{\varepsilon}}-T_{\mu_{\varepsilon},\nu_{\varepsilon}}\left(1+\mathrm{O}_{p}\left(\frac{\sigma}{n}\right)\right)\right\rangle.$ (12)
@@
$\sum_{x\in G_{n}}S_{\mu_{\varepsilon},\nu_{\varepsilon}}(x)-T_{\mu_{\varepsilon},\nu_{\varepsilon}}(x)=\mathrm{O}_{p}(\sigma n),$ (13)
@@
In general, one can hope that the ratio $\sigma/n$ is small, so that the result suggests that understanding the quantity $\sup_{f}\left\langle f,S_{\mu_{\varepsilon},\nu_{\varepsilon}}\right\rangle+\left\langle f^{\mathsf{d}},T_{\mu_{\varepsilon},\nu_{\varepsilon}}\right\rangle$ under a suitable choice of noise is a first step to take towards completely characterizing the impact of the noise.
@@ 3.2 Noise model
The previous section invites us to consider a noise model for which it is not necessary to rescale the measures. To this end, we will consider slightly correlated Gaussian noise where we identify each coordinate of the Gaussian noise vector with a point on the regular grid $G_{n}$.
@@
Note that this noise model is equivalent to drawing the pixels independently from $\mathcal{N}\left(0,\sigma^{2}m/(m-1)\right)$, calculating their mean, and then subtracting the mean from every pixel.
@@
This last property allows us to focus on the impact of the noise, while setting aside the questions pertaining to rescaling the measures whose behavior was captured in Theorem 1.